OpenStudy (anonymous):

I need to prove 1*1! + 2*2!+...+n*n!=(n+1)!-1 with mathematical induction

OpenStudy (badhi):

for n=1, RHS=(1+1)!-1=2-1=1=LHS. therefore true for n=1, assume true for n=p, $$1\times1!+2\times2!+...+p\times p!=(p+1)!-1$$ add (p+1)\times (p+1)! to both sides, $$\begin{align*} 1\times1!+2\times2!+...+p\times p!+(p+1)\times(p+1)!&=(p+1)!-1+(p+1)\times (p+1)!\\ &=(p+1)!\left[1+p+1\right]-1\\ &=(p+1)!\times (p+2)-1\\ &=(p+2)!-1\\ &=[(p+1)+1]-1 \end{align*}$$ therefore it is true for n=p+1